OpenStudy (anonymous):

Solving Special Systems Classify system. Give the number of solutions. y-2x=5 x=y-3

OpenStudy (anonymous):

@texaschic101 Can you help ?

OpenStudy (texaschic101):

There is only one solution x = y - 3 x + 3 = y or y = x + 3 y - 2x = 5 y = 2x + 5